OpenStudy (anonymous):

in probability what does E[|X|] mean? I know that |X| means cardinality but what does |X| mean when we use it in sense of probability?

OpenStudy (anonymous):

you can also find this symbol at http://en.wikipedia.org/wiki/Expected_value

OpenStudy (anonymous):

Isn't that random variable ?

OpenStudy (anonymous):

sorry I mean absolute value instead of cardinality in the original question****

OpenStudy (anonymous):

yes X is a random variable but when we use perpendicular bars | | in E[|X|] then what does it mean?

OpenStudy (anonymous):

That is the absolute value of the random variable for expectation that is E..

OpenStudy (anonymous):

OK @waterineyes Also, in martingale when we use E[X_n+1 || F_n] = X_n then the two perpendicular bars || denotes conditional probability or what?

OpenStudy (anonymous):

Conditional Probability is shown by one bar only I think.. Like this : \[P(A|B)\]

OpenStudy (anonymous):

yeah that I know but when we use || then what does it mean?

OpenStudy (anonymous):

This is sign for Absolute Value...

OpenStudy (anonymous):

Suppose : |x| = 5y + 6 So here, to eliminate bars from x: There will form two equations: |x| = 5y + 6 gives x = 5y + 6 as first equation.. |x| = 5y + 6 gives -x = 5y + 6 as second equation...

OpenStudy (anonymous):

I do understand the concept of absolute value but I don't understand the meaning of || in ecepectation

OpenStudy (anonymous):

I am explaining it in terms of math.. But in probability the concept is same..

OpenStudy (anonymous):

In that link read the column named "INEQUALITY"